The following table contains the number of complaints received in the department store for first 6 months of operation:

Month

Jan

Feb

Mar

April

May

June

Complaints

36

45

81

90

108

144

  1. a. Calculate the forecast for the month from April through June using 3 month moving average? (3 points)
  2. b. What will be the forecast for the month of July using 3 month moving average? (3 points)
  3. c. If actual sales for the month of July turned out to be 145, what will be the forecast for August month using 3 month moving Average? (4 points)
Answer #1

a. Using the 3 month moving average method

F(April) = {Actual(Jan) + Actual(Feb) + Actual(Mar)}/3

= (36+45+81)/3

= 54

F(May) = (45+81+90)/3

= 72

F(June) = (81+90+108)/3

= 93

b. F(July) = (90+108+144)/3

= 114

c. F(August) = (108+144+145)/3

= 132.33